what is hesses law please explane with suitable example

Grade:12

5 Answers

Siddharth Singh
6 years ago
 Hess' Law states that the heat evolved or absorbed in a chemical process is the same whether the process takes place in one or in several steps.
Example-
  • A + B = AB, dH1
  • AB + B = AB2dH2 
    then,
    A + 2 B = AB2dH1 2 = dH1 + dH2
Gitanjali Rout
6 years ago
If a reaction takes place in several steps then its standard reaction enthalphy is sum of standard enthalphy of those intermediate reaction into which ovarLl reaction into which may bedivided at same temperature
Gitanjali Rout
6 years ago
Reaction that take place at several steps then its standard reaction enthalphy is the sum of standard enthalphy of those
Vasanth Chavan
6 years ago
Hess's Law of Constant Heat Summation (or just Hess's Law) states that regardless of the multiple stages or steps of a reaction, the total enthalpy change for the reaction is the sum of all changes
 
Anjali kumari
5 years ago
Hess law states the change of enthalpy of rection in a chemical reaction. It's one example is Methanol.